# Learning Constrained Markov Decision Processes With Non-stationary Rewards and Constraints

### v1

@@ Appendix B Events dictionary
    $\bigg{\lvert}\widehat{g}_{t,i}(x,a)-g^{\circ}(x,a)\bigg{\rvert}\leq\xi_{t}(x,a).$
@@ Appendix C Confidence intervals
$\bigg{\lvert}{g}^{\circ}(x,a)-\frac{1}{T}\underset{t\in[T]}{\sum}\mathbb{E}[g_{t,i}(x,a)]\bigg{\rvert}\leq\frac{C_{G}}{T}.$
@@
$\bigg{|}g^{\circ}(x,a)-\frac{1}{T}\underset{t\in[T]}{\sum}\mathbb{E}[g_{t,i}(x,a)]\bigg{|}$ $=\bigg{|}\frac{1}{T}\underset{t\in[T]}{\sum}(g^{\circ}(x,a)-\mathbb{E}[g_{t,i}(x,a)])\bigg{|}$
$\leq\frac{1}{T}\underset{t\in[T]}{\sum}\bigg{|}g^{\circ}(x,a)-\mathbb{E}[g_{t,i}(x,a)]\bigg{|}$
@@
Notice that the proof holds for all $i\in[m]$ since $g^{\circ}$ and $C_{G}$ are defined employing the maximum over $i\in[m]$. Following the same steps, it holds:
@@
$\bigg{\lvert}\widehat{g}_{t,i}(x,a)-g^{\circ}(x,a)\bigg{\rvert}\leq\sqrt{\frac{1}{2\max\{N_{t}(x,a),1\}}\ln\left(\frac{2}{\delta}\right)}+\frac{C_{G}}{\max\{N_{t}(x,a),1\}}.$
@@
$\bigg{|}\widehat{g}_{t,i}(x,a)-g^{\circ}(x,a)\bigg{|}$ $=\left|\left(\frac{\sum_{\tau\in[t]}\mathbb{I}_{\tau}(x,a)g_{\tau,i}(x,a)}{\max\{N_{t}(x,a),1\}}\right)-g^{\circ}(x,a)\right|$
@@
$\ \ +\left|\frac{1}{\max\{N_{t}(x,a),1\}}\underset{\tau\in[t]}{\sum}\mathbb{I}_{\tau}(x,a)[\mathbb{E}[g_{\tau,i}(x,a)]-g^{\circ}(x,a)]\right|,$ (9)
@@
$\Bigg{|}\frac{1}{\max\{N_{t}(x,a),1\}}\underset{\tau\in[t]}{\sum}\mathbb{I}_{\tau}(x,a)$ $\left[\mathbb{E}[g_{\tau,i}(x,a)]-g^{\circ}(x,a)\right]\Bigg{|}$
$\leq\frac{1}{\max\{N_{t}(x,a),1\}}\underset{\tau\in[t]}{\sum}\mathbb{I}_{\tau}(x,a)\bigg{|}\mathbb{E}[g_{\tau,i}(x,a)]-g^{\circ}(x,a)\bigg{|}$
$\leq\frac{1}{\max\{N_{t}(x,a),1\}}\underset{\tau\in[T]}{\sum}\bigg{|}\mathbb{E}[g_{\tau,i}(x,a)]-g^{\circ}(x,a)\bigg{|}$
@@
$\bigg{\lvert}\widehat{g}_{t,i}(x,a)-g^{\circ}(x,a)\bigg{\rvert}\leq\sqrt{\frac{1}{2\max\{N_{t}(x,a),1\}}\ln\left(\frac{2mT|X||A|}{\delta}\right)}+\frac{C_{G}}{\max\{N_{t}(x,a),1\}}.$
@@
$\mathbb{P}\Bigg{[}\bigg{\lvert}\widehat{g}_{t,i}(x,a)-g^{\circ}(x,a)\bigg{\rvert}\leq\zeta_{t}(x,a)\Bigg{]}\geq 1-\delta^{\prime}.$
@@
$\mathbb{P}\Bigg{[}\bigcap_{x,a,i,t}\Big{\{}\Big{|}\widehat{g}_{t,i}(x,a)-g^{\circ}(x,a)\Big{|}\leq\zeta_{t}(x,a)\Big{\}}\Bigg{]}.$
@@
$\mathbb{P}\Bigg{[}\bigcap_{x,a,i,t}\Big{\{}\Big{|}\widehat{g}_{t,i}(x,a)-g^{\circ}(x,a)$ $\Big{|}\leq\zeta_{t}(x,a)\Big{\}}\Bigg{]}$
$=1-\mathbb{P}\Bigg{[}\bigcup_{x,a,i,t}\Big{\{}\Big{|}\widehat{g}_{t,i}(x,a)-g^{\circ}(x,a)\Big{|}\leq\zeta_{t}(x,a)\Big{\}}^{c}\Bigg{]}$
$\geq 1-\sum_{x,a,i,t}\mathbb{P}\Bigg{[}\Big{\{}\Big{|}\widehat{g}_{t,i}(x,a)-g^{\circ}(x,a)\Big{|}\leq\zeta_{t}(x,a)\Big{\}}^{c}\Bigg{]}$ (10)
@@
$\bigg{\lvert}\widehat{g}_{t,i}(x,a)-g^{\circ}(x,a)\bigg{\rvert}\leq\sqrt{\frac{1}{2\max\{N_{t}(x,a),1\}}\ln\left(\frac{2mT|X||A|}{\delta}\right)}+\frac{C_{G}}{\max\{N_{t}(x,a),1\}},$
@@
$\leq\Bigg{\lvert}\widehat{g}_{t,i}(x,a)-g^{\circ}(x,a)\Bigg{\rvert}+\Bigg{\lvert}{g}^{\circ}(x,a)-\frac{1}{T}\underset{t\in[T]}{\sum}\mathbb{E}[g_{t,i}(x,a)]\Bigg{\rvert}$
@@ Appendix D Omitted proofs when the corruption is known
$=\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left[\left(\mathbb{E}[g_{t,i}]-g^{\circ}\right)^{\top}q_{t}+g^{\circ}{}^{\top}q_{t}-\alpha_{i}\right]^{+}$
$\leq\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left[\left(\mathbb{E}[g_{t,i}]-g^{\circ}\right)^{\top}q_{t}+\left(\underline{g}_{t-1,i}+2\xi_{t-1}\right)^{\top}q_{t}-\alpha_{i}\right]^{+}$ (12a)
$=\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left[\left(\mathbb{E}[g_{t,i}]-g^{\circ}\right)^{\top}q_{t}+\underline{g}_{t-1,i}^{\top}\left(q_{t}-\widehat{q}_{t}\right)+\underline{g}_{t-1,i}^{\top}\widehat{q}_{t}+2\xi_{t-1}^{\top}q_{t}-\alpha_{i}\right]^{+}$
$\leq\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left[\left(\mathbb{E}[g_{t,i}]-g^{\circ}\right)^{\top}q_{t}+\underline{g}_{t-1,i}^{\top}\left(q_{t}-\widehat{q}_{t}\right)+2\xi_{t-1}^{\top}q_{t}\right]^{+}$ (12b)
$\leq\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left|\left(\mathbb{E}[g_{t,i}]-g^{\circ}\right)^{\top}q_{t}\right|+2\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left|\xi_{t-1}^{\top}q_{t}\right|+\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left|\underline{g}_{t-1,i}^{\top}\left(q_{t}-\widehat{q}_{t}\right)\right|$ (12c)
$\leq\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left\lVert\mathbb{E}[g_{t,i}]-g^{\circ}\right\rVert_{1}+2\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\xi_{t-1}^{\top}q_{t}+\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left\lVert q_{t}-\widehat{q}_{t}\right\rVert_{1}$ (12d)
@@
$\leq 3|X||A|\sqrt{\frac{1}{2}T\ln\left(\frac{2mT|X||A|}{\delta}\right)}+|X||A|(2+\ln(T))C_{G}+|X||A|C_{G}+L\sqrt{2T\ln\frac{1}{\delta}}$ (13c)
$\leq 3|X||A|\sqrt{\frac{1}{2}T\ln\left(\frac{2mT|X||A|}{\delta}\right)}+(3+\ln(T))|X||A|C_{G}+L\sqrt{2T\ln\frac{1}{\delta}}$
$=\mathcal{O}\left(|X||A|\sqrt{T\ln\left(\frac{mT|X||A|}{\delta}\right)}+\ln(T)|X||A|C_{G}\right),$
where Inequality (13a) follows from the Azuma-Hoeffding inequality and noticing that $\sum_{x,a}\xi_{t-1}(x,a)q_{t}(x,a)\leq L$, Equality (13b) follows from the definition of $\xi_{t}$ and finally, Inequality (13c) holds since $1+\sum_{t=1}^{N_{T}(x,a)}\sqrt{\frac{1}{t}}\leq 1+2\sqrt{N_{T}(x,a)}\leq 3\sqrt{N_{T}(x,a)}$ and $1+\sum_{t=1}^{N_{T}(x,a)}\frac{1}{t}\leq 2+\ln(T)$. Finally, we notice that the intersection event $\mathcal{E}_{G}\cap\mathcal{E}_{\widehat{q}\ }\cap\mathcal{E}_{\text{Azuma}}$ holds with the following probability,
@@
$\leq 3|X||A|\sqrt{\frac{1}{2}T\ln\left(\frac{2T|X||A|}{\delta}\right)}+|X||A|(2+\ln(T))C_{r}+|X||A|C_{r}+L\sqrt{2T\ln\frac{1}{\delta}}$ (14c)
$\leq 3|X||A|\sqrt{\frac{1}{2}T\ln\left(\frac{2T|X||A|}{\delta}\right)}+(3+\ln(T))|X||A|C_{r}+L\sqrt{2T\ln\frac{1}{\delta}}$
$=\mathcal{O}\left(|X||A|\sqrt{T\ln\left(\frac{T|X||A|}{\delta}\right)}+\ln(T)|X||A|C_{r}\right),$
where Inequality (14a) follows from Azuma-Hoeffding inequality, Equality (14b) holds for the definition of $\phi_{t}$, and Inequality (14c) holds since $1+\sum_{t=1}^{N_{T}(x,a)}\sqrt{\frac{1}{t}}\leq 1+2\sqrt{N_{T}(x,a)}\leq 3\sqrt{N_{T}(x,a)}$ and $1+\sum_{t=1}^{N_{T}(x,a)}\frac{1}{t}\leq 2+\ln(T)$. Thus, we observe that with probability at least $1-9\delta$ it holds:
$R_{T}=\mathcal{O}\left(L|X|\sqrt{|A|T\ln\left(\frac{T|X||A|}{\delta}\right)}+|X||A|\sqrt{T\ln\left(\frac{T|X||A|}{\delta}\right)}+\ln(T)|X||A|C_{r}\right).$
Employing Theorem 3 and the definition of $C$, which is at least equal to $C_{r}$, concludes the proof. ∎
@@ Appendix E Omitted proofs when the knowledge of $C$ is not precise
$g^{\circ}(x,a)\leq\widehat{g}_{t,i}(x,a)+\xi_{t}(x,a)+\left(\frac{\epsilon}{\max\{N_{t}(x,a),1\}}+\frac{\epsilon}{T}\right).$
@@
$g^{\circ}(x,a)\leq\underline{g}_{t,i}(x,a)+2\xi_{t}(x,a)+\left(\frac{\epsilon}{\max\{N_{t}(x,a),1\}}+\frac{\epsilon}{T}\right).$
@@
$\bigg{\lvert}\widehat{g}_{t,i}(x,a)-$ $g^{\circ}(x,a)]\bigg{\rvert}\leq$
@@
$\bigg{\lvert}\widehat{g}_{t,i}(x,a)-g^{\circ}(x,a)]\bigg{\rvert}\leq\xi_{t}(x,a)+\frac{\epsilon}{\max\{N_{t}(x,a),1\}}+\frac{\epsilon}{T},$
@@
$R_{T}=\mathcal{O}\left(L|X||A|\sqrt{T\ln\left(\frac{T|X||A|}{\delta}\right)}+\ln(T)|X||A|\widehat{C}\right).$
@@
$V_{T}=\mathcal{O}\left(L|X||A|\sqrt{T\ln\left(\frac{mT|X||A|}{\delta}\right)}+\ln(T)|X||A|C_{G}\right).$
@@
$=\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left[\left(\mathbb{E}[g_{t,i}]-g^{\circ}\right)^{\top}q_{t}+g^{\circ}{}^{\top}q_{t}-\alpha_{i}\right]^{+}$
$\leq\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\Bigg{[}(\mathbb{E}[g_{t,i}]-g^{\circ})^{\top}q_{t}+\underline{g}_{t-1,i}^{\top}(q_{t}-\widehat{q}_{t})+\underline{g}_{t-1,i}^{\top}\widehat{q}_{t}+2\xi_{t-1}^{\top}q_{t}+$
@@
$V_{T}=\mathcal{O}\left(L|X|\sqrt{|A|T\ln\left(\frac{T|X||A|}{\delta}\right)}+|X||A|\sqrt{T\ln\left(\frac{mT|X||A|}{\delta}\right)}+\ln(T)|X||A|C_{G}\right),$
@@
$V_{T}=\mathcal{O}\left(L|X||A|\sqrt{T\ln\left(\frac{T|X||A|}{\delta}\right)}+\ln(T)|X||A|\widehat{C}\right).$

### v2

@@ Appendix B Events dictionary
    $\bigg{\lvert}\widehat{g}_{t,i}(x,a)-g_{i}^{\circ}(x,a)\bigg{\rvert}\leq\xi_{t}(x,a),$
    where $g_{i}^{\circ}\in[0,1]^{|X\times A|}\coloneq[G^{\circ}]_{i}$.
@@ Appendix C Confidence intervals
$\bigg{\lvert}{g}_{i}^{\circ}(x,a)-\frac{1}{T}\underset{t\in[T]}{\sum}\mathbb{E}[g_{t,i}(x,a)]\bigg{\rvert}\leq\frac{C_{G}}{T}.$
@@
$\bigg{|}g^{\circ}_{i}(x,a)-\frac{1}{T}\underset{t\in[T]}{\sum}\mathbb{E}[g_{t,i}(x,a)]\bigg{|}$ $=\bigg{|}\frac{1}{T}\underset{t\in[T]}{\sum}(g^{\circ}_{i}(x,a)-\mathbb{E}[g_{t,i}(x,a)])\bigg{|}$
$\leq\frac{1}{T}\underset{t\in[T]}{\sum}\bigg{|}g^{\circ}_{i}(x,a)-\mathbb{E}[g_{t,i}(x,a)]\bigg{|}$
@@
Notice that the proof holds for all $i\in[m]$ since $C_{G}$ is defined employing the maximum over $i\in[m]$. Following the same steps, it holds:
@@
$\bigg{\lvert}\widehat{g}_{t,i}(x,a)-g^{\circ}_{i}(x,a)\bigg{\rvert}\leq\sqrt{\frac{1}{2\max\{N_{t}(x,a),1\}}\ln\left(\frac{2}{\delta}\right)}+\frac{C_{G}}{\max\{N_{t}(x,a),1\}}.$
@@
$\bigg{|}\widehat{g}_{t,i}(x,a)-g^{\circ}_{i}(x,a)\bigg{|}$ $=\left|\left(\frac{\sum_{\tau\in[t]}\mathbb{I}_{\tau}(x,a)g_{\tau,i}(x,a)}{\max\{N_{t}(x,a),1\}}\right)-g^{\circ}_{i}(x,a)\right|$
@@
$\ \ +\left|\frac{1}{\max\{N_{t}(x,a),1\}}\underset{\tau\in[t]}{\sum}\mathbb{I}_{\tau}(x,a)[\mathbb{E}[g_{\tau,i}(x,a)]-g^{\circ}_{i}(x,a)]\right|,$ (10)
@@
$\Bigg{|}\frac{1}{\max\{N_{t}(x,a),1\}}\underset{\tau\in[t]}{\sum}\mathbb{I}_{\tau}(x,a)$ $\left[\mathbb{E}[g_{\tau,i}(x,a)]-g^{\circ}_{i}(x,a)\right]\Bigg{|}$
$\leq\frac{1}{\max\{N_{t}(x,a),1\}}\underset{\tau\in[t]}{\sum}\mathbb{I}_{\tau}(x,a)\bigg{|}\mathbb{E}[g_{\tau,i}(x,a)]-g^{\circ}_{i}(x,a)\bigg{|}$
$\leq\frac{1}{\max\{N_{t}(x,a),1\}}\underset{\tau\in[T]}{\sum}\bigg{|}\mathbb{E}[g_{\tau,i}(x,a)]-g^{\circ}_{i}(x,a)\bigg{|}$
@@
$\bigg{\lvert}\widehat{g}_{t,i}(x,a)-g^{\circ}_{i}(x,a)\bigg{\rvert}\leq\sqrt{\frac{1}{2\max\{N_{t}(x,a),1\}}\ln\left(\frac{2mT|X||A|}{\delta}\right)}+\frac{C_{G}}{\max\{N_{t}(x,a),1\}}.$
@@
$\mathbb{P}\Bigg{[}\bigg{\lvert}\widehat{g}_{t,i}(x,a)-g^{\circ}_{i}(x,a)\bigg{\rvert}\leq\zeta_{t}(x,a)\Bigg{]}\geq 1-\delta^{\prime}.$
@@
$\mathbb{P}\Bigg{[}\bigcap_{x,a,i,t}\Big{\{}\Big{|}\widehat{g}_{t,i}(x,a)-g^{\circ}_{i}(x,a)\Big{|}\leq\zeta_{t}(x,a)\Big{\}}\Bigg{]}.$
@@
$\mathbb{P}\Bigg{[}\bigcap_{x,a,i,t}\Big{\{}\Big{|}\widehat{g}_{t,i}(x,a)-g^{\circ}_{i}(x,a)$ $\Big{|}\leq\zeta_{t}(x,a)\Big{\}}\Bigg{]}$
$=1-\mathbb{P}\Bigg{[}\bigcup_{x,a,i,t}\Big{\{}\Big{|}\widehat{g}_{t,i}(x,a)-g^{\circ}_{i}(x,a)\Big{|}\leq\zeta_{t}(x,a)\Big{\}}^{c}\Bigg{]}$
$\geq 1-\sum_{x,a,i,t}\mathbb{P}\Bigg{[}\Big{\{}\Big{|}\widehat{g}_{t,i}(x,a)-g^{\circ}_{i}(x,a)\Big{|}\leq\zeta_{t}(x,a)\Big{\}}^{c}\Bigg{]}$ (11)
@@
$\bigg{\lvert}\widehat{g}_{t,i}(x,a)-g^{\circ}_{i}(x,a)\bigg{\rvert}\leq\sqrt{\frac{1}{2\max\{N_{t}(x,a),1\}}\ln\left(\frac{2mT|X||A|}{\delta}\right)}+\frac{C_{G}}{\max\{N_{t}(x,a),1\}},$
@@
$\leq\Bigg{\lvert}\widehat{g}_{t,i}(x,a)-g^{\circ}_{i}(x,a)\Bigg{\rvert}+\Bigg{\lvert}{g}^{\circ}_{i}(x,a)-\frac{1}{T}\underset{t\in[T]}{\sum}\mathbb{E}[g_{t,i}(x,a)]\Bigg{\rvert}$
@@ Appendix D Omitted proofs when the corruption is known
$=\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left[\left(\mathbb{E}[g_{t,i}]-g^{\circ}_{i}\right)^{\top}q_{t}+g^{\circ}_{i}{}^{\top}q_{t}-\alpha_{i}\right]^{+}$
$\leq\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left[\left(\mathbb{E}[g_{t,i}]-g^{\circ}_{i}\right)^{\top}q_{t}+\left(\underline{g}_{t-1,i}+2\xi_{t-1}\right)^{\top}q_{t}-\alpha_{i}\right]^{+}$ (13a)
$=\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left[\left(\mathbb{E}[g_{t,i}]-g^{\circ}_{i}\right)^{\top}q_{t}+\underline{g}_{t-1,i}^{\top}\left(q_{t}-\widehat{q}_{t}\right)+\underline{g}_{t-1,i}^{\top}\widehat{q}_{t}+2\xi_{t-1}^{\top}q_{t}-\alpha_{i}\right]^{+}$
$\leq\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left[\left(\mathbb{E}[g_{t,i}]-g^{\circ}_{i}\right)^{\top}q_{t}+\underline{g}_{t-1,i}^{\top}\left(q_{t}-\widehat{q}_{t}\right)+2\xi_{t-1}^{\top}q_{t}\right]^{+}$ (13b)
$\leq\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left|\left(\mathbb{E}[g_{t,i}]-g^{\circ}_{i}\right)^{\top}q_{t}\right|+2\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left|\xi_{t-1}^{\top}q_{t}\right|+\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left|\underline{g}_{t-1,i}^{\top}\left(q_{t}-\widehat{q}_{t}\right)\right|$ (13c)
$\leq\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left\lVert\mathbb{E}[g_{t,i}]-g^{\circ}_{i}\right\rVert_{1}+2\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\xi_{t-1}^{\top}q_{t}+\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left\lVert q_{t}-\widehat{q}_{t}\right\rVert_{1}$ (13d)
@@
$\leq 3\sqrt{\frac{1}{2}|X||A|LT\ln\left(\frac{2mT|X||A|}{\delta}\right)}+|X||A|(2+\ln(T))C_{G}+|X||A|C_{G}+L\sqrt{2T\ln\frac{1}{\delta}}$ (14c)
$\leq 3\sqrt{\frac{1}{2}|X||A|LT\ln\left(\frac{2mT|X||A|}{\delta}\right)}+(3+\ln(T))|X||A|C_{G}+L\sqrt{2T\ln\frac{1}{\delta}}$
$=\mathcal{O}\left(\sqrt{|X||A|LT\ln\left(\frac{mT|X||A|}{\delta}\right)}+\ln(T)|X||A|C_{G}\right),$
where Inequality (14a) follows from the Azuma-Hoeffding inequality and noticing that $\sum_{x,a}\xi_{t-1}(x,a)q_{t}(x,a)\leq L$, Equality (14b) follows from the definition of $\xi_{t}$ and finally, Inequality (14c) holds since $1+\sum_{t=1}^{N_{T}(x,a)}\sqrt{\frac{1}{t}}\leq 1+2\sqrt{N_{T}(x,a)}\leq 3\sqrt{N_{T}(x,a)}$ , since $1+\sum_{t=1}^{N_{T}(x,a)}\frac{1}{t}\leq 2+\ln(T)$ and by Cauchy-Schwarz inequality. Finally, we notice that the intersection event $\mathcal{E}_{G}\cap\mathcal{E}_{\widehat{q}\ }\cap\mathcal{E}_{\text{Azuma}}$ holds with the following probability,
@@
$\leq 3\sqrt{\frac{1}{2}|X||A|LT\ln\left(\frac{2T|X||A|}{\delta}\right)}+|X||A|(2+\ln(T))C_{r}+|X||A|C_{r}+L\sqrt{2T\ln\frac{1}{\delta}}$ (15c)
$\leq 3\sqrt{\frac{1}{2}|X||A|LT\ln\left(\frac{2T|X||A|}{\delta}\right)}+(3+\ln(T))|X||A|C_{r}+L\sqrt{2T\ln\frac{1}{\delta}}$
$=\mathcal{O}\left(\sqrt{|X||A|LT\ln\left(\frac{T|X||A|}{\delta}\right)}+\ln(T)|X||A|C_{r}\right),$
where Inequality (15a) follows from Azuma-Hoeffding inequality, Equality (15b) holds for the definition of $\phi_{t}$, and Inequality (15c) holds since $1+\sum_{t=1}^{N_{T}(x,a)}\sqrt{\frac{1}{t}}\leq 1+2\sqrt{N_{T}(x,a)}\leq 3\sqrt{N_{T}(x,a)}$, $1+\sum_{t=1}^{N_{T}(x,a)}\frac{1}{t}\leq 2+\ln(T)$ and by Cauchy-Schwarz inequality. Thus, we observe that with probability at least $1-9\delta$ it holds:
$R_{T}=\mathcal{O}\left(L|X|\sqrt{|A|T\ln\left(\frac{T|X||A|}{\delta}\right)}+\ln(T)|X||A|C_{r}\right).$
Employing Corollary 3 and the definition of $C$, which is at least equal to $C_{r}$, concludes the proof. ∎
@@ Appendix E Omitted proofs when the knowledge of $C$ is not precise
$g^{\circ}_{i}(x,a)\leq\widehat{g}_{t,i}(x,a)+\xi_{t}(x,a)+\left(\frac{\epsilon}{\max\{N_{t}(x,a),1\}}+\frac{\epsilon}{T}\right).$
@@
$g^{\circ}_{i}(x,a)\leq\underline{g}_{t,i}(x,a)+2\xi_{t}(x,a)+\left(\frac{\epsilon}{\max\{N_{t}(x,a),1\}}+\frac{\epsilon}{T}\right).$
@@
$\bigg{\lvert}\widehat{g}_{t,i}(x,a)-$ $g^{\circ}_{i}(x,a)]\bigg{\rvert}\leq$
@@
$\bigg{\lvert}\widehat{g}_{t,i}(x,a)-g^{\circ}_{i}(x,a)]\bigg{\rvert}\leq\xi_{t}(x,a)+\frac{\epsilon}{\max\{N_{t}(x,a),1\}}+\frac{\epsilon}{T},$
@@
$R_{T}=\mathcal{O}\left(L|X|\sqrt{|A|T\ln\left(\frac{T|X||A|}{\delta}\right)}+\ln(T)|X||A|\widehat{C}\right).$
@@
$V_{T}=\mathcal{O}\left(L|X|\sqrt{|A|T\ln\left(\frac{mT|X||A|}{\delta}\right)}+\ln(T)|X||A|C_{G}\right).$
@@
$=\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\left[\left(\mathbb{E}[g_{t,i}]-g^{\circ}_{i}\right)^{\top}q_{t}+g^{\circ}_{i}{}^{\top}q_{t}-\alpha_{i}\right]^{+}$
$\leq\underset{i\in[m]}{\max}\underset{t\in[T]}{\sum}\Bigg{[}(\mathbb{E}[g_{t,i}]-g^{\circ}_{i})^{\top}q_{t}+\underline{g}_{t-1,i}^{\top}(q_{t}-\widehat{q}_{t})+\underline{g}_{t-1,i}^{\top}\widehat{q}_{t}+2\xi_{t-1}^{\top}q_{t}+$
@@
$V_{T}=\mathcal{O}\left(L|X|\sqrt{|A|T\ln\left(\frac{T|X||A|}{\delta}\right)}+\ln(T)|X||A|C_{G}\right),$
@@
$V_{T}=\mathcal{O}\left(L|X|\sqrt{|A|T\ln\left(\frac{T|X||A|}{\delta}\right)}+\ln(T)|X||A|\widehat{C}\right).$